Given the equation of the ellipse a. The Center b. The Foci c. Width of major...

  1. Given the equation of the ellipse (x-4)^2/25 +(y+2)^2/49=1

  2. a. The Center

  3. b. The Foci

  4. c. Width of major axis
Answer #1

Hey there stay home and stay safe

\small \frac{(x-4)^2}{25}+\frac{(y+2)^2}{49}=1\\ \Rightarrow \frac{(x-4)^2}{5^2}+\frac{(y+2)^2}{7^2}=1\\ \\$ Equating with $ \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1 $ we have$\\ h=4,k=-2a=5,b=7\\ $Centre is (h,k)$=(4,-2)\\ $foci are $(h,k\pm c)\\ c^2=b^2-a^2=7^2-5^2=49-25=24\Rightarrow c=2\sqrt6\\ foci=(4,-2\pm2\sqrt6)\\ Major\ axis=2b=2\times7=14

And please leave a like. Thank you.